OpenStudy (anonymous):

Solve for x in the proportion x/x+12 = 1/x-1

OpenStudy (helder_edwin):

u have \[ \large \frac{x}{x+12}=\frac{1}{x-1} \] then \[ \large x(x-1)=1(x+12) \] can u finish this?

OpenStudy (anonymous):

um, well what do i do next?

OpenStudy (helder_edwin):

it is cuadratic equation \[ \large x^2-x=x+12 \] u should be able to solve this can u?

OpenStudy (anonymous):

yeah

OpenStudy (helder_edwin):

great! do it and tell me what u get just to check

OpenStudy (anonymous):

wait, so i should use the quadratic equation?

OpenStudy (helder_edwin):

do it whichever way u want (i) quadratic formula (ii) complete the square (iii) factoring

OpenStudy (anonymous):

okay

OpenStudy (helder_edwin):

for instance \[ \large x^2-x=x+12 \] \[ \large x^2-2x+1=12+1=13 \] then \[ \large (x-1)^2=13 \] \[ \large x-1=\pm\sqrt{13} \] \[ \large x=1\pm\sqrt{13} \]

OpenStudy (helder_edwin):

this is completing the square u should get the same if you use the formula

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i got x ≈ −2.61 and x ≈ 4.61

OpenStudy (helder_edwin):

u r right

OpenStudy (anonymous):

okay thanks

OpenStudy (helder_edwin):

u r w

OpenStudy (anonymous):

could you help me with another?